OpenStudy (anonymous):

Help Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 A. x = 0 B. x = 2 C. There is no x value as there is no solution. D. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

What do you think is the solution?

OpenStudy (anonymous):

i dont know. thats why im asking

OpenStudy (anonymous):

Have you tried to work out the problem?

OpenStudy (anonymous):

i dont know how to

OpenStudy (anonymous):

@aai y u block me its me Peaches remember

OpenStudy (anonymous):

Hello Peaches

OpenStudy (flvs.net):

Still need help?

OpenStudy (anonymous):

@Nixy hey

OpenStudy (anonymous):

yeah...

OpenStudy (flvs.net):

What do you think the answer is?

OpenStudy (anonymous):

I already asked that Flvs

OpenStudy (anonymous):

need help

OpenStudy (flvs.net):

What does he think?

OpenStudy (anonymous):

He just said he does not know

OpenStudy (flvs.net):

@kitty_lover1234 are you as cute as your name is?

OpenStudy (anonymous):

Look at the top. I dont know what it is, thats why im asking. i just want to know how to do it

OpenStudy (anonymous):

what

OpenStudy (anonymous):

OMFG

OpenStudy (flvs.net):

Lol, i'm just kidding.

OpenStudy (anonymous):

MATH. PLEASE.

OpenStudy (flvs.net):

Okay aai, one sec.

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

.... what do i think the answer is?

OpenStudy (flvs.net):

2x + y = 1 4x + 2y = −1 rearrange the first like this: y = 1 -2x now sub in this as y in the second: 4x + 2(1-2x) = -1 4x - 4x + 2 = -1 2 = -1 hence inconsistent. no solutions

OpenStudy (anonymous):

i agree

OpenStudy (flvs.net):

Thanks @kitty_lover1234

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

so its C?

OpenStudy (anonymous):

It is C

OpenStudy (flvs.net):

Yep, That's why they call me the Human Calculator. :D

OpenStudy (anonymous):

i think so

OpenStudy (flvs.net):

A medal would be nice.....

OpenStudy (anonymous):

lol

OpenStudy (flvs.net):

I mean I just did your work, lol

OpenStudy (anonymous):

@kitty_lover1234 its true u have ta give him a medal for helping

OpenStudy (flvs.net):

@Peaches15 I help @aai

OpenStudy (anonymous):

•_• thank you

OpenStudy (flvs.net):

Youe welcome, anymore?

OpenStudy (anonymous):

sorry, i was distracted

OpenStudy (anonymous):

gave u your meadel

OpenStudy (flvs.net):

Oh thanks.

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

thank you for the meadel

OpenStudy (flvs.net):

No Problemo

OpenStudy (flvs.net):